OpenStudy (me123):

Multiply -7/2 (-5/8)

OpenStudy (anonymous):

to multiple fractions you multiply the numerators together, and then do the same for the denominators. -7 times -5 is 35, which is your new numerator. 2 times 8 is 16, your new denominator. the answer is 35/16.

OpenStudy (anonymous):

To multiply two fractions you just multiply the numerator, and multiply the denominator. \[\frac{a}{b} * \frac{c}{d} = \frac{a*c}{b*d} \]

OpenStudy (anonymous):

-7/2 (-5/8)=(-7*-5)/(2*8) = 35/16

OpenStudy (me123):

Thank you